OpenStudy (anonymous):

Substitute your birthday into the equation Square root of the quantity x minus y + m = d, where y is the last two digits of your birth year, m is the month, and d is the day. If you were born on 7/10/1856 like Nikola Tesla, your equation would be Square root of the quantity x minus 56 + 7 = 10. Solve for x and identify if it is an extraneous solution. my7 bday is 3/8/1996

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@helpme1.2

OpenStudy (anonymous):

@radar

OpenStudy (whpalmer4):

I've got an easy one: same month and same day, \(x = \text{year}\) :-)

OpenStudy (whpalmer4):

@radar are you doing your birth year? you must be either much younger or older than you look :-) DI_KAZ' birth year is 96, not 93.

OpenStudy (radar):

My year is 38 (1938) lol

OpenStudy (whpalmer4):

if you were born a/b/cd, the problem is \[\sqrt{x-cd} + a = b\]

OpenStudy (anonymous):

i got the qution \[\sqrt{x-96}+3=8\]

OpenStudy (whpalmer4):

based on the Tesla example

OpenStudy (whpalmer4):

Right, so how to solve \[\sqrt{x-96}+3=8\]I'd start by moving the 3 to the right hand side. Then you can square both sides and solve for \(x\).

OpenStudy (anonymous):

can u show me the steps im a visaul learner

OpenStudy (whpalmer4):

Okay, subtract 3 from both sides of the equation. What's the new equation?

OpenStudy (anonymous):

from which sides the 93 and the 8?

OpenStudy (anonymous):

96*

OpenStudy (whpalmer4):

\[\sqrt{x-96} + 3 = 8\]Subtract 3 from both sides of the equation: \[\sqrt{x-96} + 3 - 3 = 8 -3\]\[\sqrt{x-96}=5\] Agreed? If we add or subtract the same quantity from both sides, the equation remains equivalent.

OpenStudy (anonymous):

RIGHT

OpenStudy (whpalmer4):

Okay, so to get rid of that pesky radical sign on the way to getting \(x=\) all alone, let's square both sides. They are equal, so they should be equal to each other after we square both sides, too. \[(\sqrt{x-96})^2 = 5^2\]But the square of the square root is just the value under the square root sign: \[x-96 = 5^2 \]\[x-96=25\] You can solve that for \(x\), right?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

here the ^2 come from

OpenStudy (whpalmer4):

^2 came from squaring each side: I could also have written \[(\sqrt{x-96})(\sqrt{x-96}) = 5*5\]\[x-96 = 25\]

OpenStudy (anonymous):

ok then what

OpenStudy (whpalmer4):

Okay, you've got some money in your pocket. If you hand me 15 bucks, you'll have 10 left. How much money do you have in your pocket?

OpenStudy (whpalmer4):

How much money would you need to have to be able to give me $15 and still have $10 for yourself?

OpenStudy (anonymous):

5

OpenStudy (whpalmer4):

If you had $5 in your wallet, you could give me $15 and have $10 left for yourself? I don't think so!

OpenStudy (anonymous):

25

OpenStudy (whpalmer4):

If you had $25 in your wallet, you could give me $15, and you would have $25-$15 = $10 left. That's exactly the form of the problem we are trying to solve. We've got some number, \(x\), and if we subtract 96 from it, we have 25 left.

OpenStudy (whpalmer4):

\[x-96=25\]If we add 96 to both sides:\[x-96+96 = 25+96\]\[x + 0 = 25 + 96\]\[x=25+96\]\[x =\]

OpenStudy (anonymous):

121

OpenStudy (whpalmer4):

yes!

OpenStudy (anonymous):

haha thanks

OpenStudy (whpalmer4):

Now we have to test our answer in the original problem to see if it works, or if it is an extraneous solution (read: it doesn't work) \[\sqrt{121-96} +3 = 8\]\[\sqrt{25}+3=8\]What is the square root of 25?

OpenStudy (whpalmer4):

what number when multiplied by itself gives you 25 as the answer?

OpenStudy (anonymous):

5

OpenStudy (whpalmer4):

Right. \[5+3=8\]True or false?

OpenStudy (radar):

@DI_KAZ Does the DI stand for Drill Instructor??

OpenStudy (anonymous):

true and yes i want to be one it helps me keep my goals in sight if i see it alot

OpenStudy (radar):

Good goal.

OpenStudy (anonymous):

thanks

OpenStudy (radar):

You're welcome. whpalmer4 really laid that problem out. Good luck with your studies.

OpenStudy (anonymous):

thanks and thanks @whpalmer4

OpenStudy (whpalmer4):

My kid brother was a Marine Harrier pilot during the first Gulf War. Semper Fi!

OpenStudy (anonymous):

that awesome !!!

OpenStudy (whpalmer4):

tag me if you ever get stuck on a problem, and I'll try to help you out when I'm next available.